I am having problems setting this up. Please help me.
An airplane travels at a specific air speed between two cities. On a certain day the airplane takes two hours to make the trip with the help of a 10 mph tailwind. Returning home later in the day, the trip takes five hours with a 30 mph headwind. Find the speed of the plane in still air.

Answer by mananth(16946)

plane speed =x mph
wind speed 10 mph

time with wind 2 hours
time against wind(30mph) 4 hours
Distance = same
2(x+10)=4(x-30)
2x+20 = 4 x -120
4x-2x=20-120
x = 140 mph plane speed